# MENTOR: Mixture-of-Experts Network with Task-Oriented Perturbation for Visual Reinforcement Learning

### v1

@@ Appendix B Simulation Experimental Settings
Appendix C Real-World Experimental Settings
The training and testing videos are available at mentor. The hyperparameters for the real-world experiments are the same as those used in the simulator, as shown in Table 2. We use 16 experts, with the top 4 experts activated.
C.1 Observation Space
@@ C.3 Reward Design
$R_{\text{peg}}=\frac{1}{2}\left(\left(\sqrt{2}-\|\mathbf{x}_{g}-\mathbf{x}_{c}\|\right)\cdot C_{1}+\left(2-|\Delta z|\right)\cdot C_{2}+\left(\frac{\pi}{2}-|\theta_{c}-\theta_{g}|\right)\cdot C_{3}-C_{4}\right)$ (6)
@@
$R_{cable}=r_{\text{slot}_{1}}+\mathbb{I}(r_{\text{slot}_{1}}\geq 2)\cdot\left(r_{\text{slot}_{2}}+C_{5}\right)$ (7)
@@
$\begin{split}R_{\mathrm{golf}}=\left(2-\|\mathbf{p}_{\text{club}}-\mathbf{p}_{\text{ball}}\|\right)\cdot C_{8}+\left(2-\|\mathbf{p}_{\text{ball}}-\mathbf{p}_{\text{hole}}\|\right)\cdot C_{9}-\mathbb{I}(\text{strike})\\ +\left(2-|\theta_{\text{best}}-\theta_{\text{current}}|\right)\cdot C_{10}-\max\left(0,\mathbf{p}_{\text{ball}}[y]-\mathbf{p}_{\text{club}}[y]+0.05\right)\cdot C_{11}\end{split}$ (8)
Where:
  • $\mathbf{p}_{\text{hole}}$: Position of the target hole.
@@
C.4 Auto-Reset Mechanisms
@@
Appendix D Time Efficiency of MENTOR
@@
Table 4: Comparison of time efficiency in the simulation task (FPS).
Task Name MENTOR DrM DrQ-v2 ALIX TACO
Hopper 37 55 78 49 23
@@
Appendix E MENTOR in Real-World Multi-Tasking Process

### v3

@@ Appendix C Ablation Study on Key Contributions
Appendix D Real-World Experimental Settings
The training and testing videos are available at mentor-vrl. The hyperparameters for the real-world experiments are the same as those used in the simulator, as shown in Table 3. We use 16 experts, with the top 4 experts activated.
D.1 Observation Space
@@ D.3 Reward Design
$R_{\text{peg}}=$ $\frac{1}{2}\bigg{(}\left(\sqrt{2}-\|\mathbf{x}_{g}-\mathbf{x}_{c}\|\right)\cdot C_{1}+\left(2-|\Delta z|\right)\cdot C_{2}$
$+\left(\frac{\pi}{2}-|\theta_{c}-\theta_{g}|\right)\cdot C_{3}-C_{4}\bigg{)}$
@@
$R_{cable}=$ $r_{\text{slot}_{1}}+\mathbb{I}(r_{\text{slot}_{1}}\geq 2)\cdot\left(r_{\text{slot}_{2}}+C_{5}\right)$
@@
$R_{\mathrm{golf}}=$ $\left(2-\|\mathbf{p}_{\text{club}}-\mathbf{p}_{\text{ball}}\|\right)\cdot C_{8}+\left(2-\|\mathbf{p}_{\text{ball}}-\mathbf{p}_{\text{hole}}\|\right)\cdot C_{9}$
$-\mathbb{I}(\text{strike})+\left(2-|\theta_{\text{best}}-\theta_{\text{current}}|\right)\cdot C_{10}$
$-\max\left(0,\mathbf{p}_{\text{ball}}[y]-\mathbf{p}_{\text{club}}[y]+0.05\right)\cdot C_{11}$
@@
Table 5: Comparison of time efficiency in the simulation task (FPS).
Task Name MENTOR DrM DrQ-v2 ALIX TACO
Hopper 37 55 78 49 23
Where:
  • $\mathbf{p}_{\text{hole}}$: Position of the target hole.
D.4 Auto-Reset Mechanisms
@@
Figure 12: Multi-task performance of MoE on Meta-World Simulator. Evaluation accuracy during training for 3, 4, 5, and 7-task combinations, highlighting MoE’s advantage over MLP in multitask settings.
Appendix E Time Efficiency of MENTOR
@@
We further extend the training process of the DrM baseline to reach the same performance level as MENTOR in three real-world experiments with the training time comparison shown in Figure 14, which demonstrates an average 37% improvement in time efficiency for our method. These findings underscore the importance of each component in achieving superior results.
Figure 13: Cosine similarity of multistage in a single task. This figure shows the cosine similarities of gradients on the corresponding four stages (Grasp, Move, Assemble, and Release) for both MLP and MoE agents.
Appendix F Mixture-of-Experts Alleviation Gradient Conflicts in Single Task
In Meta-World, manipulation tasks are associated with compound reward functions that typically include components such as reaching, grasping, and placing. Conflicts between these objectives can arise, creating a burden for shared parameters.
To validate this, we analyze the gradient cosine similarities for the Assembly task. The Assembly task, as shown in Figure 4 can naturally be divided into four stages: Grasp, Move, Assemble, and Release.
To illustrate how Mixture-of-Experts alleviates gradient conflicts in a single task, we evaluate the cosine similarities of gradients on the corresponding four stages for both MLP and MoE agents, as shown in Figure 13. The result show that the MLP agent experiences gradient conflicts between grasping and the other stages. This can occur because the procedure of reaching to grasp objects could increase the distance between the robot and the target pillar, leading to competing optimization signals. In contrast, the MoE agent mitigates these conflicts, achieving consistently positive gradient cosine similarities across all stage pairs. This validates the ability of the MoE architecture to alleviate the burden of shared parameters and facilitate more efficient optimization, even in single-task scenarios.
Appendix G MENTOR in Simulator Multi-Tasking Process
To further demonstrate the multi-task capabilities of MoE, we conducted additional multitask learning experiments on Meta-World. In these experiments, we used four task combinations, consisting of 3, 4, 5, and 7 tasks, respectively. Figure 12 shows the evaluation accuracy during training and the detailed composition of the multi-tasks. In this experiment, neither MENTOR nor DrM used perturbation, with the only difference being the use of MoE versus MLP, indicating the effectiveness of MoE in multi-tasking process.
Appendix H MENTOR in Real-World Multi-Tasking Process
@@
Appendix I Random Disturbances in Simulation
To demonstrate the generalization capabilities of the agents trained by MENTOR, we have introduced random disturbances in the real-world experiments presented in Section 4.2. Additionally, we make evaluation of Meta-World Assembly task with random disturbance. In detail, the training phase remains unchanged, but during evaluation, we introduce a random disturbance: after the robot grasps the ring and moves toward the fitting area, the fitting pillar randomly changes its location (Disturbance). This forces the robot agent to adjust its trajectory to the new target position. Figure 16 shows the agent consistently accomplishes Assembly task even with the disturbance, showing the policies learned by MENTOR exhibit strong robustness against disturbances.
Figure 16: Random disturbances in simulation. This figure shows the execution of the learned agent using MENTOR. The agent consistently accomplishes Assembly task even with the disturbance.